OpenStudy (anonymous):

Write a point-slope equation for the line with the given slope that contains the given point... m= -3; ( -2,-5)

OpenStudy (aroub):

y+5=-3(x+2)

OpenStudy (anonymous):

How did you get that answer ?

OpenStudy (aroub):

u use the point slope form : y-y1=m(x-x1) and u substitute -2 by y1 and -5 by x1 and -3 by m

OpenStudy (anonymous):

Okay Thank you so much